### 16 Mar 2026 v2

# A characterization of Fano type varieties

Abstract.
We prove a characterization of Fano type varieties.
2020 Mathematics Subject Classification:
14E30
1. Introduction
Let $X$ be a normal variety. We say that $X$ is of klt type if there is an effective $\mathbb{Q}$-divisor $\Delta$ such that $(X,\Delta)$ is klt. Let $X$ be a projective normal variety. We say that $X$ is of Fano type if there is an effective $\mathbb{Q}$-divisor $\Delta$ such that $(X,\Delta)$ is klt and $(K_{X}+\Delta)$ is ample.
In [zhuang2024direct], Zhuang proved that pure images of klt type singularities are of klt type. The following klt type criterion is crucial in his proof.
Lemma $1.$
[zhuang2024direct, Lemma 2.4] Let $X$ be a normal variety. Then $X$ is of klt type if and only if
  1. (1)
    the anticanonical ring $R(X,-K_{X}):=\bigoplus_{m\geq 0}O_{X}(-mK_{X})$ is finitely generated over $O_{X}$, and
  2. (2)
    $X^{\prime}:=\operatorname{Proj}_{X}R(X,-K_{X})$ is klt.
In this note, we prove the following theorem, which is a global version of $1.$
Theorem 1.2.
Let $X$ be a projective normal variety. Then $X$ is of Fano type if and only if
  1. (1)
    $K_{X}$ is big,
  2. (2)
    the anticanonical ring $R(X,-K_{X}):=\bigoplus_{m\geq 0}H^{0}(X,-mK_{X})$ is finitely generated, and
  3. (3)
    $Y:=\operatorname{Proj}_{\mathbb{C}}R(X,-K_{X})$ is klt.
Varieties of Fano type are not necessarily $\mathbb{Q}$-Gorenstein. Hence Theorem 1.2 generalizes [CasciniGongyo2013anticanonical, Theorem 1.1] or [ChoiHwangPark2015Factorization, Theorem 1.1], where the $\mathbb{Q}$-Gorenstein assumption on the variety is required.
We discuss section rings for Weil divisors in Section 2, and prove Theorem 1.2 in Section 3.
Acknowledgments
The author thanks Professor Chen Jiang for providing a simpler proof of Theorem 1.2. The author acknowledges Professor Mao Sheng for the financial support through the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-032).
Notation and conventions
We work over the field of complex numbers.
Let $f:X\to Y$ be a projective morphism with connected fibers between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. There are effective $\mathbb{Q}$-divisors $D^{h}$ and $D^{v}$ such that $D=D^{h}+D^{v}$ and a component $P$ of $D$ dominates $Y$ if and only if $P\subset\operatorname{Supp}D^{h}$. We call such $D^{h}$ and $D^{v}$ the horizontal part and the vertical part of $D$, respectively. We say that $D$ is vertical over $Y$ if $D=D^{v}$. If $D$ is vertical over $Y$, we say that $D$ is exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $f(Q)=P$ but $Q\not\subset\operatorname{Supp}D$.
For a Weil divisor $D$ on a projective normal variety $X$, we denote by $f_{D}$ the rational map associated to the linear system $|D|$. For a linear system $|D|$, we denote by $\operatorname{Mov}|D|$ and $\operatorname{Fix}|D|$ its moving part and fixed part, respectively.
2. On Weil divisors whose section rings are finitely generated
In this section, unless otherwise specified, a divisor refers to a Weil divisor.
Setting.
Let $X$ be a projective normal variety and let $D$ be a (non-necessarily Cartier) divisor on $X$, such that $R(X,D):=\bigoplus_{m\geq 0}H^{0}(X,mD)$ is finitely generated and is generated in degree one. Let $Y:=\operatorname{Proj}R(X,D)$. Let $f=f_{D}:X\dashrightarrow\mathbb{P}^{N}$ be the rational map associated to the linear system $|D|$. Let $\pi:X^{\prime}\to X$ be a resolution of singularities, such that the rational map $f^{\prime}:X^{\prime}\dashrightarrow Y$ is a morphism. Let $\operatorname{Ex}(\pi)$ be the codimension one part of the exceptional locus of $\pi$. ∎
This section aims to generalize the following well-known theorem to Weil divisors.
Theorem 2.1.
In the Setting. Assume that $D$ is Cartier. Then $Y$ is normal, $Y=f^{\prime}(X^{\prime})$, $f^{\prime}$ has connected fibers, the rational maps $f_{mD}$ are birationally equivalent to $f$ for all $m>0$, and there is an ample and free Cartier divisor AA on $Y$, such that $\operatorname{Mov}|m\pi^{*}D|=|mf^{\prime*}A|$ for all $m>0$.
We start with the following Lemma.
Lemma $2.$
In the Setting.
  1. (1)
    There is a divisor $D^{\prime}$ on $X^{\prime}$ such that
[…]
  2. (2)
    Among all divisors satisfying 2.0.1, there is a minimal one.
  3. (3)
    If $D^{\prime}$ is a divisor satisfying 2.0.1 and $E$ is an effective $\pi$-exceptional divisor, then $D^{\prime}+E$ also satisfies 2.0.1.
  4. (4)
    If $D^{\prime}$ is a divisor satisfying 2.0.1, then $\pi_{*}:H^{0}(mD^{\prime})\cong H^{0}(mD)$ for all $m\geq 0$.
Proof.
(1) Let $\{s_{i}\}$ be basis of $H^{0}(D)\subset\mathbb{C}(X)$. Then a divisor $D^{\prime}$ with $\pi_{*}D^{\prime}=D$ satisfies 2.0.1 if and only if $\operatorname{div}_{X^{\prime}}(s_{i})+D^{\prime}\geq 0$ for all ii. Hence $\pi^{-1}_{*}D+m\operatorname{Ex}(\pi)$ satisfies 2.0.1 for all $m\gg 0$.
(2) If $D_{1}$, $D_{2}$ are divisors satisfying 2.0.1, then $D_{1}\wedge D_{2}$ also satisfies 2.0.1.
(3) It is clear from the argument in (1).
(4) It follows from the commutative diagram
[…]
Theorem 2.3.
In the Setting. We have $Y$ is normal, $Y=f^{\prime}(X^{\prime})$, $f^{\prime}$ has connected fibers, the rational maps $f_{mD}$ are birationally equivalent to $f$ for all $m>0$, and the following holds.
  1. (1)
    There is an ample and free Cartier divisor AA on $Y$ such that if $D^{\prime}$ is a divisor satisfying 2.0.1, then $\operatorname{Mov}|mD^{\prime}|=|mf^{\prime*}A|$ for all $m>0$.
  2. (2)
    If $D^{\prime}$ is a divisor satisfying 2.0.1, then the $f^{\prime}$-vertical part of $\operatorname{Fix}|D^{\prime}|$ is very exceptional over $Y$.
  3. (3)
    The $f^{\prime}$-vertical part of $\operatorname{Ex}(\pi)$ is very exceptional over $Y$.
Proof.
By Lemma 2.2 and Theorem 2.1, all statements except (1), (2), and (3) are clear.
Let $\tilde{D}$ be the minimal one among all divisors satisfying 2.0.1. Write $|\tilde{D}|=|M|+\tilde{F}$ as a sum of moving part and fixed part. By Lemma 2.2 and Theorem 2.1, there is an ample and free Cartier divisor AA on $Y$ such that $\operatorname{Mov}|m\tilde{D}|=|mf^{\prime*}A|$ and $\operatorname{Fix}|m\tilde{D}|=m\tilde{F}$ for all $m>0$.
We prove (1) and (2) in this paragraph. Let $D^{\prime}$ be a divisor satisfying 2.0.1. Since $\tilde{D}$ is minimal, there is an effective divisor $E$ such that $D^{\prime}=\tilde{D}+E$. By Lemma 2.2(4), we have $|mD^{\prime}|=|m\tilde{D}|+mE$. Hence $\operatorname{Mov}|mD^{\prime}|=\operatorname{Mov}|m\tilde{D}|=|f^{\prime*}mA|$. Let $F^{\prime}$ be the $f^{\prime}$-vertical part of $\operatorname{Fix}|D^{\prime}|$. By $\operatorname{Fix}|m(M+F^{\prime})|=mF^{\prime}$ for all $m>0$ and [birkar2012existence, Lemma 3.2], we have $F^{\prime}$ is very exceptional over $Y$.
(3) Since $\tilde{D}+\operatorname{Ex}(\pi)$ satisfies 2.0.1 and $\operatorname{Ex}(\pi)\leq\operatorname{Fix}|\tilde{D}+\operatorname{Ex}(\pi)|$, we have the $f^{\prime}$-vertical part of $\operatorname{Ex}(\pi)$ is very exceptional over $Y$ by (2). ∎
3. Proof of Theorem 1.2
Lemma 3.1.
Let $(X,\Delta)$ be a projective klt pair with $(K_{X}+\Delta)$ nef and big, then $X$ is of Fano type.
Proof.
Since $(K_{X}+\Delta)$ is big, there are an ample $\mathbb{Q}$-divisor AA and an effective $\mathbb{Q}$-divisor $E$ such that $(K_{X}+\Delta)\sim_{\mathbb{Q}}A+E$. If $0<\epsilon\ll 1$, then $(X,\Delta+\epsilon E)$ is klt and $(K_{X}+\Delta+\epsilon E)\sim_{\mathbb{Q}}(1-\epsilon)(A+E)+\epsilon A$ is ample. ∎
Lemma 3.2.
Let $f:X\to Y$ be a birational morphism between projective normal varieties. If $X$ is of Fano type, then $Y$ is of Fano type.
Proof.
Let $\Delta$ and AA be effective $\mathbb{Q}$-divisors such that AA is ample, $(X,\Delta+A)$ is klt and $K_{X}+\Delta+A\sim_{\mathbb{Q}}0$. By $K_{X}+\Delta+A\sim_{\mathbb{Q}}f^{*}(K_{Y}+f_{*}\Delta+f_{*}A)$, we have $(Y,f_{*}\Delta+f_{*}A)$ is klt. Since $f_{*}A$ is big, there are an ample effective $\mathbb{Q}$-divisor $A^{\prime}$ and an effective $\mathbb{Q}$-divisor $E$ such that $f_{*}A\sim_{\mathbb{Q}}A^{\prime}+E$. If $0<\epsilon\ll 1$, then $(Y,f_{*}\Delta+(1-\epsilon)f_{*}A+\epsilon E)$ is klt and $(K_{Y}+f_{*}\Delta+(1-\epsilon)f_{*}A+\epsilon E)\sim_{\mathbb{Q}}\epsilon A$ is ample. ∎
Proof of Theorem 1.2.
The “only if” part. Since $X$ is of Fano type, we have $K_{X}$ is big and $X$ is of klt type. By Zhuang’s Lemma 1.1, we have $\bigoplus_{m\geq 0}O_{X}(-mK_{X})$ is finitely generated over $O_{X}$. Consider the birational morphism $\pi:X^{\prime}:=\operatorname{Proj}_{X}\bigoplus_{m\geq 0}O_{X}(-mK_{X})\to X$. By [KM98, Lemma 6.2], $\pi$ is small, $K_{X^{\prime}}$ is $\mathbb{Q}$-Cartier, and
(3.0.1) $\pi_{*}O_{X^{\prime}}(-mK_{X^{\prime}})=O_{X}(-mK_{X})\text{ for all }m\geq 0.$
Hence $X^{\prime}$ is a $\mathbb{Q}$-Gorenstein Fano type variety by [Birkar2016FanotypeFibraiton, Lemma 2.4]. By [CasciniGongyo2013anticanonical, Theorem 1.1] and 3.0.1, we have $R(X,-K_{X})\cong R(X^{\prime},-K_{X^{\prime}})$ is finitely generated and $Y=\operatorname{Proj}R(X,-K_{X})\cong\operatorname{Proj}R(X^{\prime},-K_{X^{\prime}})$ is klt.
The “if” part. Let $r$ be a positive integer such that $R(X,-rK_{X})$ is generated in degree one. Let $f:X\dashrightarrow\mathbb{P}^{N}$ be the rational map associated to the linear system $|-rK_{X}|$. Let $\pi:X^{\prime}\to X$ be a resolution of singularities such that the rational map $f^{\prime}:X^{\prime}\dashrightarrow\mathbb{P}^{N}$ is a morphism. Let $\operatorname{Ex}(\pi)$ be the codimension one part of the exceptional locus of $\pi$. By Lemma 2.2, there is an effective and $\pi$-exceptional divisor $E$ on $X^{\prime}$ such that
$\pi_{*}:|m(-rK_{X^{\prime}}+E)|\to|-mrK_{X}|\text{ is one-one for all }m>0.$
Write $|-rK_{X^{\prime}}+E|=|M|+F$ as a sum of moving part and fixed part. By Theorem 2.3, we have
  1. (1)
    $Y=f^{\prime}(X^{\prime})$,
  2. (2)
    $f^{\prime}$ has connected fibers,
  3. (3)
    $M=f^{\prime*}A$ for an ample Cartier divisor AA on $Y$, and
  4. (4)
    the $f^{\prime}$-vertical part of $F$ and $\operatorname{Ex}(\pi)$ are very exceptional over $Y$.
Since $K_{X}$ is big and $f^{\prime}$ has connected fibers, we have $f^{\prime}$ is a birational morphism. Hence $F$ and $\operatorname{Ex}(\pi)$ are exceptional over $Y$. Hence, the rational map $f:X\dashrightarrow Y$ is a birational contraction, i.e., there are no $f^{-1}$-exceptional divisors on $Y$. By
$rK_{Y}=f^{\prime}_{*}(-rK_{X^{\prime}})=f^{\prime}_{*}(-rK_{X^{\prime}}+E)=f^{\prime}_{*}(M+F)=f^{\prime}_{*}(f^{\prime*}A+F)=A,$
we have
(3.0.2) $K_{X^{\prime}}-\frac{1}{r}E+\frac{1}{r}F\sim_{\mathbb{Q}}f^{*}K_{Y}.$
Let $P\subset X^{\prime}$ be any prime divisor which is $f^{\prime}$-exceptional and non $\pi$-exceptional. By 3.0.2, we have
$a(P,Y)=a(P,X^{\prime},-\frac{1}{r}E+\frac{1}{r}F)=-\operatorname{mult}_{P}(-\frac{1}{r}E+\frac{1}{r}F)=-\operatorname{mult}_{P}\frac{1}{r}F\leq 0.$
By [BCHM, Corollary 1.4.3], there is a projective klt pair $(Z,\Delta_{Z})$ and a birational morphism $g:Z\to Y$ such that $K_{Z}+\Delta_{Z}\sim_{\mathbb{Q}}g^{*}K_{Y}$, and the induced rational map $Z\dashrightarrow X$ is small. Since $K_{Y}$ is ample, we have $(K_{Z}+\Delta_{Z})$ is nef and big. Hence $Z$ is of Fano type by Lemma 3.1. By [Birkar2016FanotypeFibraiton, Lemma 2.4], $X$ is of Fano type.
Another Proof of the “if” part (Provided by Chen Jiang). By [BCHM], there is a projective $\mathbb{Q}$-factorial terminal variety $Y$ and a birational morphism $\pi:X^{\prime}\to X$ such that $K_{X^{\prime}}$ is nef over $X$.
We claim that $\pi_{*}:|-mK_{X^{\prime}}|\to|-mK_{X}|$ is one-one for all $m\geq 0$. Let $s$ be a rational function on $X$ such that $\operatorname{div}_{X}(s)-mK_{X}\geq 0$. Then $\operatorname{div}_{X^{\prime}}(s)-mK_{X^{\prime}}$ is a $\pi$-anti-nef divisor with $\pi_{*}(\operatorname{div}_{X^{\prime}}(s)-mK_{X^{\prime}})=\operatorname{div}_{X}(s)-mK_{X}\geq 0$. By the Negativity Lemma, we have $\operatorname{div}_{X^{\prime}}(s)-mK_{X^{\prime}}\geq 0$. Hence, the claim follows.
Since $X^{\prime}$} is $\mathbb{Q}$-Gorenstein, $R(Y,-K_{X^{\prime}})\cong R(X,-K_{X})$ is finitely generated and $\operatorname{Proj}R(X^{\prime},-K_{X^{\prime}})\cong\operatorname{Proj}R(X,-K_{X})$ is klt, we have $X^{\prime}$ is of Fano type by [CasciniGongyo2013anticanonical, Theorem 1.1]. By Lemma 3.2, $X$ is of Fano type. ∎